(b) If events A and B are such that they are independent, and P(A) = 0.3 with P(B) = 0.5;


i. Find P(A n B) and P(AUB)


ii. Are A and B mutually exclusive? Explain.



1
Expert's answer
2021-12-24T09:01:56-0500

i.


"P(A\cap B)=P(A)P(B)=0.3(0.5)=0.15"

"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=0.3+0.5-0.15=0.65"

ii.


"P(A\cup B)=0.65\not=0.3+0.5=P(A)+P(B)"

A and B are not mutually exclusive.